# Model Editing by Pure Fine-Tuning

### 16 Feb 2024 v1

@@ 4 Related Work
Table 4: Fluency and consistency with mass-editing on CounterFact.
@@ 5.1 Mass-Editing
Our main results are on mass-editing with on ZsRE and CounterFact, following the same setting in Meng et al. (2023). Each dataset provides 10,000 requested edits (i.e., $\mathcal{E}$). As described in Section 3, we augment the fine-tuning data with pseudo-paraphrases of the edit prompts $\mathcal{P}$ for generalization supervision and random facts $\mathcal{R}$ for locality supervision. More specifically, for each edit in $\mathcal{E}$ we generate 15 paraphrases and take 20 facts from the training split while ensuring that they do not contain any evaluation facts. Thus the total number of facts we fine-tune on is 360,000. We fine-tune GPT-J (6B) (Wang and Komatsuzaki, 2021) with LoRA (Hu et al., 2022) for computational efficiency. We optimize (2). The training takes around 7 hours on 8 GPUs.
@@ 5.3 Generative metrics
We report results on generative metrics in Table 4 and 5 for completeness. Following prior work (Meng et al., 2022, 2023), we report fluency (entropy of $n$-gram distributions) and consistency (similarity score with a reference text). Our fine-tuning methods take a hit on these metrics while improving the edit score, showing that more work is needed to go beyond classification. However, recent work shows that none of the compared editors preserves downstream performance (Gu et al., 2024), thus achieving this goal meaningfully is still an open problem.
@@ 6 Conclusions
We have demonstrated that pure fine-tuning is sufficient to obtain strong edit performance with a slight modification: optimizing the conditional likelihood and augmenting the data with additional facts. Our results challenge the assumption that fine-tuning is ineffective as a model editor, suggesting that model editing could become part of standard training.

### 10 Mar 2024 v2

@@ 3 Method
Editor Score Efficacy Generalization Locality
— (GPT-J) 37.4 34.4 (1.7) 34.5 (1.5) 45.3 (0.9)
ROME 35.0 39.8 (2.2) 25.5 (1.4) 46.9 (0.8)
MEMIT 67.3 99.2 (0.3) 80.2 (1.2) 45.3 (0.8)
FT 55.8 68.1 (1.8) 60.4 (1.7) 44.5 (0.8)
FT + M + P + R 68.5 99.6 (0.2) 84.6 (1.1) 45.8 (0.9)
Table 3: Mass-editing results on WikiRecent (1,266 edits).
@@ 4 Related Work
     + Wikipedia Loss 84.8 609.2 (0.6) 29.2 (0.2)
Table 5: Fluency and consistency with mass-editing on CounterFact. “+ Wikipedia Loss” means we additionally optimize a language modeling loss on Wikipedia text (Section 5.3).
@@ 5.1 Mass-Editing
Our main results are on mass-editing with on ZsRE and CounterFact, following the same setting in Meng et al. (2023). Each dataset provides 10,000 requested edits (i.e., $\mathcal{E}$). As described in Section 3, we augment the fine-tuning data with pseudo-paraphrases of the edit prompts $\mathcal{P}$ for generalization supervision and random facts $\mathcal{R}$ for locality supervision. More specifically, for each edit in $\mathcal{E}$ we generate 15 paraphrases and take 20 facts from the training split while ensuring that they do not contain any evaluation facts. Thus the total number of facts we fine-tune on is 360,000. We fine-tune GPT-J (6B) (Wang and Komatsuzaki, 2021) with LoRA (Hu et al., 2022) for computational efficiency. We optimize (2). The training takes around 2-3 hours on 8 GPUs. [^1] The code is available at: https://github.com/au-revoir/model-editing-ft.
@@
5.1.1 Additional experiments on WikiRecent
To further validate our mass-editing results, we perform additional experiments on a simplified version of the WikiRecent dataset (Cohen et al., 2023). The edit requests are 1,266 recent facts in Wikidata; details are provided in Appendix A. Table 3 shows the results. We again find that while vanilla fine-tuning is not effective, our conditional fine-tuning with random data augmentation is competitive with MEMIT.
5.2 Single-Editing
@@ 5.3 Generative Metrics
We report results on generative metrics in Table 5 and 6 for completeness. Following prior work (Meng et al., 2022, 2023), we report fluency (entropy of $n$-gram distributions) and consistency (similarity score with a reference text). Our fine-tuning methods take a hit on these metrics while improving the edit score, showing that more work is needed to go beyond classification. However, recent work shows that none of the compared editors preserves downstream performance (Gu et al., 2024), thus achieving this goal meaningfully is still an open problem.
We additionally show that we can easily incorporate considerations for generative performance into fine-tuning. Specifically, we optimize $(1-\gamma)L_{1}(\theta)+\gamma L_{2}(\theta)$ where $L_{1}(\theta)$ is the loss in (2) and $L_{2}(\theta)=-\sum_{w\in\mathcal{W}}\;\log p_{\theta}(w)$ is a language modeling loss on random Wikipedia text. We choose Wikipedia articles so that they do not overlap with those used for consistency evaluation. We set $\gamma=0.1$. With this change (“+ Wikipedia Loss”), we obtain a significant improvement in both fluency and consistency at the cost of a modest drop in the edit score.
@@ 6 Conclusions
We have demonstrated that pure fine-tuning is sufficient to obtain strong edit performance with a slight modification: optimizing the conditional likelihood and augmenting the data with additional facts. Our results challenge the assumption that pure fine-tuning is ineffective as a model editor, suggesting that model editing could become part of standard training.
@@
Appendix A WikiRecent Dataset
For additional mass-editing experiments, we create a simplified version of the WikiRecent dataset (Cohen et al., 2023). The original dataset consists of 570 and 1,266 facts from Wikidata for training and testing, where each fact has been recently modified (after July 2022). An example prompt is $\pi=$“The name of the position held by Nicolaus Bergius is” with the target object $o=$“bishop”. Since the dataset does not provide paraphrase prompts for evaluating generalization, we create paraphrases of $\pi$ by few-shot prompting GPT (gpt-3.5-turbo-0125). Since the task is relatively straightforward, we find the generated paraphrases preserve the original meaning (e.g., “The position held by Nicolaus Bergius is”). For neighborhood prompts, we select 10 random neighborhood prompts from CounterFact from the corresponding train/test splits. The training data for our fine-tuning method consists of 1,266 edit requests, 18,990 augmented paraphrases (i.e., randomly generated), and 12,660 random neighborhood prompts from CounterFact (not overlapping with those in the test set).